OpenStudy (anonymous):

[x+1]=x to the 2nd power- 5

OpenStudy (anonymous):

\[x+1={^{x2}}-5\]

OpenStudy (anonymous):

do you want 2 solve it?

OpenStudy (anonymous):

yea but idk how to

OpenStudy (anonymous):

collect like terms first,factorize then solve: x+1=x^2-5 x^2-x-6=0 (x-3)(x+2)=0 x=3 or x=-2

OpenStudy (anonymous):

ok